OpenStudy (anonymous):

Find the slope through (3,5) , parallel to y=5x-1

OpenStudy (anonymous):

y-y1=m(x-x1) plug in

OpenStudy (anonymous):

y-5=5(x-3)

OpenStudy (anonymous):

simplify y-5=5x-15 y=5x-10

OpenStudy (anonymous):

parallel lines always have the same slope - 5x

OpenStudy (anonymous):

so your answer is just 5 as slope. i dont get the point of doing the problem since you already know the slope is 5